1. Recall from Example 1 that whenever Suzan sees a bag of marbles, she grabs a handful at random. She has seen a bag containing three red marbles, four green ones, three white ones, and two purple ones. She grabs seven of them. Find the probability of the following event, expressing it as a fraction in lowest terms. HINT [See Example 1.]

She has all the red ones.

2.  Recall from Example 1 that whenever Suzan sees a bag of marbles, she grabs a handful at random. She has seen a bag containing four red marbles, four green ones, two white ones, and three purple ones. She grabs eight of them. Find the probability of the following event, expressing it as a fraction in lowest terms. HINT [See Example 1.]

She has at least one green one.

3.  Recall from Example 1 that whenever Suzan sees a bag of marbles, she grabs a handful at random. She has seen a bag containing four red marbles, three green ones, two white ones, and two purple ones. She grabs five of them. Find the probability of the following event, expressing it as a fraction in lowest terms. HINT [See Example 1.]

She has two green ones and one of each of the other colors.
Answer #1

1. We have total 13 marbles here out of which 4 are Red.

She grabs 8 of them.

She can grab all 4 red and any other 4 in: 4C4*9C4 ways

She can grab any 8 in : 13C8 ways

So required Probability = 0.0979

2. We have total 13 marbles here out of which 5 are green.

She picks 7 at random

Let X= event of grabbing atleast one green.

So we need, P(X>=1) = 1- P(X=0) = 1- 5C0*8C7/13C7 = 0.99534

4. We have total 10 marbles out of which 3 are red, 3 are green, 2 are white and 2 are purple.

She oucks 5 marbles.

We want her to pick 2 green marbles and one of each other colors.

This can be done is: 3C2*3C1*2C1*2C1 ways = 36

She can pick any 5 in 10C5 ways = 252

So Required Probability = 36/252= 0.14286

.

i hope it helps..

If you have any doubts please comment and please don't dislike.

PLEASE GIVE ME A LIKE. ITS VERY IMPORTANT FOR ME.